nolegirl24:

PLEASE HELP!!!!!!!!!! There are 30 homes in Neighborhood A. Each year, the number of homes increases by 20%. Just down the road, Neighborhood B has 45 homes. Each year, 3 new homes are built in Neighborhood B. Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points) Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points) Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. (4 points)

nolegirl24:

@Falconmaster

katherine:

Part A: Write functions to represent the number of homes in Neighborhood A and Neighborhood B throughout the years. (4 points) \(y = 30 (1.2)^x\) \(y= 45 + 3*x\) Part B: How many homes does Neighborhood A have after 5 years? How many does Neighborhood B have after the same number of years? (2 points) \(30(1.2)^5= 74.65\) \(45 + 3*5= 60\) Part C: After approximately how many years is the number of homes in Neighborhood A and Neighborhood B the same? Justify your answer mathematically. solve \(30 (1.2)^x = 45 + 3*x\) using a graphing calculator i got x = \(3.32079\) in the third year the number of homes are the same

nolegirl24:

THANK YOU!!!!!!!!!!!!!!!!!

katherine:

You're welcome